Suppose a tank contains 849 m3 of neon (Ne) at an absolute pressure of 1.01×105 Pa. The temperature is changed from 293.2 to 295.9 K. What is the increase in the internal energy of the neon?
Answer #1

Given tank contains neon gas

V = 849 m^3

P = 1.01×10^5 pa

Ti = 293.2 K

Tf = 295.9 K

by using ideal gas law

n = PV/RT = (849)(1.05×10^5) / (8.314)(293.2)

= 36569.83

We know internal energy is

U = (3/2)*n*R*T

= (3/2)*(36569.83)(8.314)(295.9-293.2)

= 1.231×10^6 J